Suppose that the demand curve for garbanzo beans

is given by

Q ¼ 20 - P. where Q is thousands of pounds of beans bought per

week and P is the price in dollars per pound.

a. How many beans will be bought at P ¼ 0?

b. At what price does the quantity demanded of

beans become 0?

c. Calculate total expenditures (P Æ Q) for beans

of each whole dollar price between the prices

identified in part a and part b.

d. What price for beans yields the highest total

expenditures?

e. Suppose the demand for beans shifted to

Q ¼ 40 -2P. How would your answers to

part a through part d change? Explain the differences intuitively and with a graph


1
Expert's answer
2021-10-26T09:33:45-0400

Solution:

a.). Q = 20 – P

When P = 0

Q = 20 – 0 = 20

Q = 20,000 pounds of beans

 

b.). Q = 20 – P

0 = 20 – P

P = 20

Price when Q is zero = $20

c.). Total expenditures = P "\times" Q = 20 "\times" 20,000 = $400,000

 

d.). The price of beans that yields the highest total expenditures = 20

 

e.). Q = 40 – 2P

When P = 0

Q = 40 – 0 = 40

Q = 40,000 pounds of beans

 

Q = 40 – 2P

0 = 40 – 2P

P = 20

Price when Q is zero = $20

Total expenditures = P "\times" Q = 20 "\times" 40,000 = $800,000

 

The price of beans that yields the highest total expenditures = 20


This is depicted by the below graph: